OpenStudy (anonymous):

help... standard form. The line through (1, -4) and parallel to 2x +3y = 4.

hartnn (hartnn):

can you find slope of 2x+3y=4 ?

OpenStudy (anonymous):

i don't know how, that is my problem....

hartnn (hartnn):

can you isolate y from 2x+3y =4 ?

OpenStudy (anonymous):

divide 3 on both sides?????

hartnn (hartnn):

start by subtracting 2x from both sides then divide 3

OpenStudy (anonymous):

y=4-2x/3

hartnn (hartnn):

compare that with y=mx+c where m= slope, so what u get 'm' as?

OpenStudy (anonymous):

4????

hartnn (hartnn):

basically the co-efficient of 'x' is slope here it is m = -2/3

hartnn (hartnn):

so, now we have slope m=-2/3 , point x1=1, y1=-4

hartnn (hartnn):

just plug these in y-y1=m (x-x1)

OpenStudy (anonymous):

y+4=-2/3(x-1)

hartnn (hartnn):

yes! just simplify that start by multiplying 3 on both sides

OpenStudy (anonymous):

3(y+4)=-2(x-1) 3y+12=-2x+2

hartnn (hartnn):

add 2x on both sides...

OpenStudy (anonymous):

2x+3y+12=-2x+2 2x+3y+12=2 2x+3y=-10

hartnn (hartnn):

correct! :)